OpenStudy (laddiusmaximus):

find the derivative of x^(2) +3x -5 using the definition of the derivative (i.e.) using limits

OpenStudy (anonymous):

i guess my explanation was not good, but if you have a question i would be happy to answer it

OpenStudy (laddiusmaximus):

satellite73 your explanation made sense except for one part.

OpenStudy (laddiusmaximus):

why does (x + h)^(2) become x^(2) 2xh?

OpenStudy (anonymous):

i hope that is not what i wrote

OpenStudy (laddiusmaximus):

(x+h)^(2)-3(x+h) -5 becomes what?

OpenStudy (anonymous):

\[(x+h)^2=(x+h)(x+h)=x^2+2xh+h^2\]

OpenStudy (laddiusmaximus):

crap!!! I didnt think to look at it like that!!

OpenStudy (laddiusmaximus):

FOIL

OpenStudy (anonymous):

and so \[(x+h)^2+3(x+h)-5=x^2+2xh+h^2+3x+3h-5\] by the distributive law

OpenStudy (laddiusmaximus):

i was thinking x^(2) + h^(2)

OpenStudy (anonymous):

it is the distributive law foil is a trick math teachers play on students

OpenStudy (laddiusmaximus):

well thats not cool

OpenStudy (laddiusmaximus):

one more thing. why is it +3(x+h) and not -3(x+h)?

OpenStudy (anonymous):

because you wrote \[f(x)=x^2+3x-5\] that is why it is + and not -

OpenStudy (laddiusmaximus):

i guess im too literal when I try to plug it into f(x) (x+h)-f(x)/h